OpenStudy (anonymous):

Suppose the area of a circle is 36π. What is its circumference?

OpenStudy (anonymous):

Area = pi(r^2) So, r = 6 in this case. Circumference = 2*pi*r = 12pi

Directrix (directrix):

12 pi is correct @GT